OpenStudy (anonymous):

which choice represents the fraction below as a single exponential expression? 6^8/6^8 A. 6^0 B. 6^16 C. 6^-16 D.6^1

jhonyy9 (jhonyy9):

6^8 ---- = 6^(8-8) = 6^0 = 1 6^8 so A. is right sure